# Antonio Ramírez-de-Arellano Dept. of Computer Science and Artificial Intelligence

### v1

@@ 4.1 Finite sets
$NVM_{F}(h_{*},i_{*},nvh_{*},wc_{*},outd_{*},\alpha^{\hbar}_{*},\alpha^{\iota}_{0})\subseteq NFIN.$
@@
The following family of sets are equal to NFIN:
@@
    $NVM_{F}(h_{1},i_{*},nvh_{*},wc_{s},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $s,t\geq 1$.
@@
    $NVM_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $r,t\geq 1$.
@@
$NVM_{F}(h_{*},i_{*},nvh_{*},wc_{*},outd_{*},\alpha^{\hbar}_{0},\alpha^{\iota}_{*})\subseteq NFIN$
@@
The following family of sets are equal to NFIN:
@@
    $NVM_{F}(h_{1},i_{*},nvh_{*},wc_{s},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $s,t\geq 1$.
@@
    $NVM_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $r,t\geq 1$.
@@
    $NVM_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{0}),$ for each $p,r,s,t,\\ u\geq 2$.
@@ 4.2 Semilinear sets
$SLIN=NVM_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{v})$
@@
$SLIN=NVM_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{*})$
@@
superset-of-or-equals This part of the proof is equal to the technique applied by the authors in [14]. That is simulating the right-linear grammar, this is possible due to the amount of viruses of each host at each moment of the computation is bounded, thus the number of possible configurations is finite. We still are in the same conditions so the same proof can be made.
For simplicity, this part of the proof has been divided in two Lemmas. Applying Lemma 4.14, the arithmetic progressions are generated by this family of VMs. Lastly, Lemma 4.16 shows the closure under union. Thus, the inclusion is formally proved.
@@
For each $n,r\geq 1,$ we have the following inclusion $\{n\cdot i+r\ |\ i\geq 1\}\in NVM_{F}(h_{2},i_{3(n+r)},nvh_{2},wc_{2},outd_{2},\alpha^{\hbar}_{2},\alpha^{\iota}_{*})$. More precisely, it will be generated by the virus machine $\Pi_{arith}$ of degree $(2,3n+3r)$ defined as:
@@
The idea of this invariant is taking the non-deterministic decision at instruction $i_{3n}$, that is, to go back to instruction $i_{1}$. In particular, $\varphi(m-1)$ is true. From this, a instruction $i_{3n}$ is reached, and we take the decision to go to instruction $i_{3n+1}$ with the following configuration $C_{3m}=(1,0,i_{3n+1},m\cdot n)$. After that, it is straightforward to reach the halting configuration.
@@
Let $Q_{1},Q_{2},\dots,Q_{m}\subseteq\mathbb{N}$ with $m>0$ arithmetical progressions, then $\cup_{j=1}^{m}Q_{j}\in NVM_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{*})$ for each $p,r,s,t,u\geq 2$.
@@ 4.3 Universality
$NRE=NVM_{F}(h_{*},i_{*},nvh_{*},wc_{s},out_{*},\alpha^{\hbar}_{*},\alpha^{\iota}_{*})$ for each $s\geq 2$.
@@ 4.4 Discussion of new ingredients
Here, table 3 summarises the previous results in addition to the new results obtained in this work with the proposed new ingredients. First, the singleton sets were characterised in the previous section with the old ingredients; note that including the new ingredients is straightforward. With finite sets, it is interesting to note that the new characterisations were previously inclusions, just fixing the host graph and/or the instruction graph to a tree. Showing the strict inclusion with the loops of size two in the host graph and size three in the instruciton graph. These loops restrictions are the same for characterising the semilinear sets, but unbounding the number of hosts instead, note that despite the $\beta$ ingredient seems to be very restrictive, we can characterise this family of sets. The most interesting result of this section arises when we unbound the size of the loops in the instruction graph and $\beta=F$, which allow us to reduce the number of necessary hosts to two. Last but not least, the universality result from [14] has been revised with the new ingredients, showing that there is a huge frontier in computational power, from finite sets with weight $1$, to $NRE$ with weight $2$.
@@ 5 Conclusions
The results in normal forms from the present work can better inform the design of applications or implementations of VMs. For instance, perhaps with ideas from [15, 16], sequels of the simulator from [19], or applications of VMs such as edge detection [20], or cryptography [21] can be improved by applying simplifications based on normal forms. Similarly, knowing which types of instruction, host, or channel graphs are used, the reachability of one configuration from another and other properties, can be decidable: in this case, searching for efficient algorithms and applications to formal verification are open problems. The results of the classes below $NRE$ help inform the computational complexity of problem solving with a model. In the case of VMs, a future direction is to investigate deeper into “sub-$NRE$” classes: in this way a better view of efficiency or lack of it, can be given for VMs. Lastly, new and optimal lower bounds are expected to improve the results of the present work, summarised in Table 3. New normal forms can perhaps include new ingredients or semantics not previously considered, such as a deeper focus on deterministic versus nondeterministic computations.

### v2

@@ 4.1 Finite sets
$\mathrm{NVM}_{F}(h_{*},i_{*},nvh_{*},wc_{*},outd_{*},\alpha^{\hbar}_{*},\alpha^{\iota}_{0})\subseteq\mathrm{NFIN}.$
@@
The following family of sets are equal to $\mathrm{NFIN}$:
@@
    $\mathrm{NVM}_{F}(h_{1},i_{*},nvh_{*},wc_{s},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $s,t\geq 1$.
@@
    $\mathrm{NVM}_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $r,t\geq 1$.
@@
$\mathrm{NVM}_{F}(h_{*},i_{*},nvh_{*},wc_{*},outd_{*},\alpha^{\hbar}_{0},\alpha^{\iota}_{*})\subseteq\mathrm{NFIN}$
@@
The following family of sets are equal to $\mathrm{NFIN}$:
@@
    $\mathrm{NVM}_{F}(h_{1},i_{*},nvh_{*},wc_{s},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $s,t\geq 1$.
@@
    $\mathrm{NVM}_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{0},\alpha^{\iota}_{0}),$ for each $r,t\geq 1$.
@@
    $\mathrm{NVM}_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{0}),$ for each $p,r,s,t,\\ u\geq 2$.
@@ 4.2 Semilinear sets
$\mathrm{SLIN}=\mathrm{NVM}_{T}(h_{*},i_{*},nvh_{r},wc_{*},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{v})$
@@
$\mathrm{SLIN}=\mathrm{NVM}_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{*})$
@@
superset-of-or-equals This part of the proof is equal to the technique applied by the authors in [14]. That is, simulating the right-linear grammar, this is possible due to the amount of viruses of each host at each moment of the computation being bounded, and thus the number of possible configurations is finite. We still are in the same conditions, so the same proof can be made.
For simplicity, this part of the proof has been divided into two lemmas. Applying Lemma 4.14, the arithmetic progressions are generated by this family of VMs. Lastly, Lemma 4.16 shows the closure under union. Thus, the inclusion is formally proved.
@@
For each $n,r\geq 1,$ we have the following inclusion $\{n\cdot i+r\ |\ i\geq 1\}\in\mathrm{NVM}_{F}(h_{2},i_{3(n+r)},nvh_{2},wc_{2},outd_{2},\alpha^{\hbar}_{2},\alpha^{\iota}_{*})$. More precisely, it will be generated by the virus machine $\Pi_{arith}$ of degree $(2,3n+3r)$ defined as:
@@
The idea of this invariant is taking the non-deterministic decision at instruction $i_{3n}$, that is, going back to instruction $i_{1}$. In particular, $\varphi(m-1)$ is true. From this, a instruction $i_{3n}$ is reached, and we take the decision to go to instruction $i_{3n+1}$ with the following configuration $C_{3m}=(1,0,i_{3n+1},m\cdot n)$. After that, it is straightforward to reach the halting configuration.
@@
Let $Q_{1},Q_{2},\dots,Q_{m}\subseteq\mathbb{N}$ with $m>0$ arithmetical progressions, then $\cup_{j=1}^{m}Q_{j}\in\mathrm{NVM}_{F}(h_{p},i_{*},nvh_{r},wc_{s},outd_{t},\alpha^{\hbar}_{u},\alpha^{\iota}_{*})$ for each $p,r,s,t,u\geq 2$.
@@ 4.3 Universality
$\mathrm{NRE}=\mathrm{NVM}_{F}(h_{*},i_{*},nvh_{*},wc_{s},out_{*},\alpha^{\hbar}_{*},\alpha^{\iota}_{*})$ for each $s\geq 2$.
@@ 4.4 Discussion of new ingredients
Here, table 3 summarises the previous results in addition to the new results obtained in this work with the proposed new ingredients. First, the singleton sets were characterised in the previous section with the old ingredients; note that including the new ingredients is straightforward. With finite sets, it is interesting to note that the new characterisations were previously inclusions, just fixing the host graph and/or the instruction graph to a tree. Showing the strict inclusion with the loops of size two in the host graph and size three in the instruciton graph. These loop restrictions are the same for characterising the semilinear sets, but unbounding the number of hosts instead, note that despite the $\beta$ ingredient seems to be very restrictive, we can characterise this family of sets. The most interesting result of this section arises when we unbound the size of the loops in the instruction graph and $\beta=F$, which allow us to reduce the number of necessary hosts to two. Last but not least, the universality result from [14] has been revised with the new ingredients, showing that there is a huge frontier in computational power, from finite sets with weight $1$, to $\mathrm{NRE}$ with weight $2$.
@@ 5 Conclusions
The results in normal forms from the present work can better inform the design of applications or implementations of VMs. For instance, perhaps with ideas from [15, 16], sequels of the simulator from [19], or applications of VMs such as edge detection [20], or cryptography [21] can be improved by applying simplifications based on normal forms. Similarly, knowing which types of instruction, host, or channel graphs are used, the reachability of one configuration from another and other properties, can be decidable: in this case, searching for efficient algorithms and applications to formal verification are open problems. The results of the classes below $\mathrm{NRE}$ help inform the computational complexity of problem solving with a model. In the case of VMs, a future direction is to investigate deeper into “sub-$NRE$” classes: in this way a better view of efficiency or lack of it, can be given for VMs. Lastly, new and optimal lower bounds are expected to improve the results of the present work, summarised in Table 3. New normal forms can perhaps include new ingredients or semantics not previously considered, such as a deeper focus on deterministic versus nondeterministic computations.